# Masked Language Modeling Becomes Conditional Density Estimation for Tabular Data Synthesis

### v1

@@ A.5.2 Evaluation Settings for Q1, Q2 and Q3
9:     $\hat{Q}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{Q}_{m}^{(s)}$
10:     $\hat{U}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{U}_{m}^{(s)}+(M+1)/M\cdot\frac{1}{M-1}\sum_{m=1}^{M}(\hat{U}_{m}^{(s)}-\bar{U}^{(s)})^{2}$,
11:where $\bar{U}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{U}_{m}^{(s)}$
12:Bias: $\frac{1}{S}\sum_{s=1}^{S}|\hat{Q}^{(s)}-Q^{*}|$
13:Coverage: $\frac{1}{S}\sum_{s=1}^{S}\mathbb{I}(Q^{*}\in(\hat{Q}^{(s)}\pm 1.96\cdot\sqrt{\hat{U}^{(s)}}))$
14:Confidence interval length: $\frac{1}{S}\sum_{s=1}^{S}2\cdot 1.96\cdot\sqrt{\hat{U}^{(s)}}$
A.6 Additional Experiments
A.6.1 Privacy Preservability
@@
  • Evaluation metrics: The $k$-anonymity property [46] is a measure used to assess the level of privacy protection in synthetic data. It ensures that each individual’s information in the dataset cannot be distinguished from that of at least $k-1$ other individuals. In other words, each record in the dataset has at least $k-1$ similar records in terms of quasi-identifiers, attributes that could potentially identify a subject. A higher $k$ value implies a higher level of anonymity and better privacy preservation.
@@
    DCR (Distance to Closest Record) [39, 56] is defined as the distances between all real training and synthetic samples. A higher DCR value indicates more effective privacy preservation, indicating a lack of overlap between the real training data and the synthetic samples. Conversely, an excessively large DCR score suggests a lower quality of the generated synthetic dataset. Therefore, the DCR metric provides insights into both the privacy-preserving capability and the quality of the synthetic dataset.
@@
    Attribute disclosure [7, 35] refers to the situation where attackers can uncover additional covariates of a record by leveraging a subset of covariates they already possess, along with similar records from the synthetic dataset. To quantify the extent to which attackers can accurately identify these additional covariates, we employ classification metrics. Higher attribute disclosure metrics indicate an increased risk of privacy leakage, implying that attackers can precisely infer unknown variables. In terms of privacy concerns, attribute disclosure can be considered a more significant issue than membership inference attacks, as attackers are assumed to have access to only a subset of covariates for a given record.
@@
  • Evaluation procedure: We evaluate the $k$-anonymity following the approach described in [40]  [^13] https://github.com/vanderschaarlab/synthcity. Additionally, similar to [56], we define DCR as the $5^{th}$ percentile of the $L_{2}$ distances between all real training samples and synthetic samples. Since DCR relies on $L_{2}$ distance and continuous variables, it is computed solely using continuous variables. We assess attribute disclosure using the methodology outlined in [7].
@@
Figure 4: Trade-off between quality and privacy. Left: feature selection performance (synthetic data quality). Right: DCR (privacy preservability). The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. Error bars represent the standard errors of the mean. The figure is identical to Figure 2.
Table 8: Trade-offs between privacy preservability and synthetic data quality (statistical fidelity and machine learning utility). The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
@@ A.6.2 Sensitivity Analysis
A.7 Detailed Experimental Results
A.7.1 Q1. Synthetic Data Quality
@@ A.7 Detailed Experimental Results
A.7.2 Q1. Visualization of Marginal Histogram
@@
A.7.3 Q2: Synthetic Data Quality in Scenarios with Incomplete Training Dataset
@@
A.7.4 Q3: Multiple Imputation Performance
Table 13: Q3: Multiple imputation under MCAR, MAR, MNARL, and MNARQ at 0.3 missingness. The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. $\downarrow$ denotes the lower is better.
@@ A.7.4 Q3: Multiple Imputation Performance
MICE $0.015_{\pm 0.001}$ $0.641_{\pm 0.033}$ $0.036_{\pm 0.002}$
GAIN $0.032_{\pm 0.002}$ $0.462_{\pm 0.034}$ $0.036_{\pm 0.002}$
VAEAC $0.013_{\pm 0.001}$ $0.698_{\pm 0.030}$ $0.036_{\pm 0.002}$
MIWAE $0.011_{\pm 0.001}$ $0.767_{\pm 0.032}$ $0.040_{\pm 0.002}$
not-MIWAE $0.016_{\pm 0.012}$ $0.693_{\pm 0.329}$ $0.050_{\pm 0.056}$
MaCoDE(MCAR) $0.014_{\pm 0.001}$ $0.753_{\pm 0.033}$ $0.041_{\pm 0.002}$
@@
MICE $0.011_{\pm 0.001}$ $0.750_{\pm 0.023}$ $0.036_{\pm 0.002}$
GAIN $0.026_{\pm 0.002}$ $0.574_{\pm 0.024}$ $0.036_{\pm 0.002}$
VAEAC $0.008_{\pm 0.000}$ $0.785_{\pm 0.022}$ $0.036_{\pm 0.002}$
MIWAE $0.007_{\pm 0.000}$ $0.865_{\pm 0.018}$ $0.040_{\pm 0.002}$
not-MIWAE $0.008_{\pm 0.001}$ $0.844_{\pm 0.020}$ $0.038_{\pm 0.002}$
MaCoDE(MAR) $0.010_{\pm 0.001}$ $0.826_{\pm 0.023}$ $0.038_{\pm 0.002}$
@@
MICE $0.016_{\pm 0.001}$ $0.611_{\pm 0.033}$ $0.036_{\pm 0.002}$
GAIN $0.035_{\pm 0.002}$ $0.414_{\pm 0.032}$ $0.036_{\pm 0.002}$
VAEAC $0.012_{\pm 0.001}$ $0.689_{\pm 0.031}$ $0.036_{\pm 0.002}$
MIWAE $0.012_{\pm 0.001}$ $0.764_{\pm 0.028}$ $0.040_{\pm 0.002}$
not-MIWAE $0.012_{\pm 0.001}$ $0.741_{\pm 0.029}$ $0.040_{\pm 0.002}$
MaCoDE(MNARL) $0.015_{\pm 0.001}$ $0.738_{\pm 0.033}$ $0.040_{\pm 0.002}$
@@
MICE $0.009_{\pm 0.000}$ $0.762_{\pm 0.023}$ $0.035_{\pm 0.002}$
GAIN $0.017_{\pm 0.001}$ $0.690_{\pm 0.025}$ $0.036_{\pm 0.002}$
VAEAC $0.008_{\pm 0.000}$ $0.796_{\pm 0.021}$ $0.036_{\pm 0.002}$
MIWAE $0.006_{\pm 0.000}$ $0.870_{\pm 0.018}$ $0.038_{\pm 0.002}$
not-MIWAE $0.007_{\pm 0.001}$ $0.852_{\pm 0.019}$ $0.038_{\pm 0.002}$
MaCoDE(MNARQ) $0.009_{\pm 0.001}$ $0.828_{\pm 0.021}$ $0.038_{\pm 0.002}$
Table 14: Q3: Multiple imputation for each dataset under MCAR at 0.3 missingness. The means and standard errors of the mean across 10 repeated experiments are reported. Due to computational issues, the number of multiple imputations is set to 10 for the covtype dataset, while for other datasets it is set to 100. $\downarrow$ denotes lower is better.
@@
Dataset concrete
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.011_{\pm 0.001}$ $0.988_{\pm 0.013}$ $0.061_{\pm 0.000}$
GAIN $0.028_{\pm 0.003}$ $0.688_{\pm 0.057}$ $0.061_{\pm 0.000}$
missMDA $0.014_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.066_{\pm 0.000}$
VAEAC $0.011_{\pm 0.001}$ $0.975_{\pm 0.017}$ $0.062_{\pm 0.000}$
MIWAE $0.007_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.065_{\pm 0.000}$
not-MIWAE $0.009_{\pm 0.001}$ $0.988_{\pm 0.012}$ $0.066_{\pm 0.000}$
EGC $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
MaCoDE $0.010_{\pm 0.001}$ $0.963_{\pm 0.019}$ $0.064_{\pm 0.000}$
Dataset covtype
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.012_{\pm 0.000}$ $0.050_{\pm 0.017}$ $0.003_{\pm 0.000}$
GAIN $0.048_{\pm 0.006}$ $0.030_{\pm 0.021}$ $0.003_{\pm 0.000}$
missMDA $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
VAEAC $0.006_{\pm 0.000}$ $0.220_{\pm 0.033}$ $0.003_{\pm 0.000}$
MIWAE $0.017_{\pm 0.001}$ $0.080_{\pm 0.029}$ $0.003_{\pm 0.000}$
not-MIWAE $0.025_{\pm 0.000}$ $0.010_{\pm 0.010}$ $0.003_{\pm 0.000}$
EGC $0.007_{\pm 0.000}$ $0.920_{\pm 0.029}$ $0.036_{\pm 0.000}$
MaCoDE $0.011_{\pm 0.000}$ $0.080_{\pm 0.029}$ $0.003_{\pm 0.000}$
Dataset kings
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.028_{\pm 0.001}$ $0.291_{\pm 0.023}$ $0.012_{\pm 0.000}$
GAIN $0.048_{\pm 0.004}$ $0.136_{\pm 0.031}$ $0.012_{\pm 0.000}$
missMDA $0.048_{\pm 0.000}$ $0.018_{\pm 0.012}$ $0.014_{\pm 0.000}$
VAEAC $0.016_{\pm 0.001}$ $0.582_{\pm 0.015}$ $0.012_{\pm 0.000}$
MIWAE $0.009_{\pm 0.001}$ $0.518_{\pm 0.043}$ $0.013_{\pm 0.000}$
not-MIWAE $0.014_{\pm 0.002}$ $0.373_{\pm 0.029}$ $0.018_{\pm 0.005}$
EGC $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
MaCoDE $0.006_{\pm 0.001}$ $0.755_{\pm 0.019}$ $0.013_{\pm 0.000}$
Dataset letter
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.024_{\pm 0.000}$ $0.219_{\pm 0.010}$ $0.014_{\pm 0.000}$
GAIN $0.054_{\pm 0.005}$ $0.131_{\pm 0.022}$ $0.013_{\pm 0.000}$
missMDA $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
VAEAC $0.024_{\pm 0.002}$ $0.281_{\pm 0.010}$ $0.014_{\pm 0.000}$
MIWAE $0.033_{\pm 0.004}$ $0.406_{\pm 0.047}$ $0.029_{\pm 0.007}$
not-MIWAE $0.032_{\pm 0.003}$ $0.550_{\pm 0.031}$ $0.026_{\pm 0.005}$
EGC $0.010_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.095_{\pm 0.001}$
MaCoDE $0.051_{\pm 0.003}$ $0.200_{\pm 0.038}$ $0.033_{\pm 0.009}$
Dataset loan
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.015_{\pm 0.001}$ $0.680_{\pm 0.044}$ $0.027_{\pm 0.000}$
GAIN $0.040_{\pm 0.006}$ $0.280_{\pm 0.074}$ $0.028_{\pm 0.000}$
missMDA $0.024_{\pm 0.000}$ $0.420_{\pm 0.020}$ $0.032_{\pm 0.000}$
VAEAC $0.022_{\pm 0.002}$ $0.420_{\pm 0.063}$ $0.028_{\pm 0.000}$
MIWAE $0.007_{\pm 0.001}$ $0.920_{\pm 0.033}$ $0.031_{\pm 0.000}$
not-MIWAE $0.011_{\pm 0.002}$ $0.840_{\pm 0.027}$ $0.032_{\pm 0.001}$
EGC $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
MaCoDE $0.023_{\pm 0.004}$ $0.780_{\pm 0.036}$ $0.030_{\pm 0.000}$
@@
Table 15: Q3: Multiple imputation for each dataset under MAR at 0.3 missingness. The means and standard errors of the mean across 10 repeated experiments are reported. Due to computational issues, the number of multiple imputations is set to 10 for the covtype dataset, while for other datasets it is set to 100. $\downarrow$ denotes lower is better.
@@
Dataset concrete
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.008_{\pm 0.001}$ $0.988_{\pm 0.013}$ $0.061_{\pm 0.000}$
GAIN $0.023_{\pm 0.002}$ $0.675_{\pm 0.057}$ $0.061_{\pm 0.000}$
missMDA $0.012_{\pm 0.001}$ $0.975_{\pm 0.017}$ $0.066_{\pm 0.000}$
VAEAC $0.007_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.061_{\pm 0.000}$
MIWAE $0.006_{\pm 0.001}$ $0.988_{\pm 0.012}$ $0.064_{\pm 0.000}$
not-MIWAE $0.006_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.064_{\pm 0.000}$
EGC $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
MaCoDE $0.009_{\pm 0.001}$ $0.963_{\pm 0.019}$ $0.063_{\pm 0.000}$
Dataset covtype
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.009_{\pm 0.001}$ $0.450_{\pm 0.027}$ $0.003_{\pm 0.000}$
GAIN $0.034_{\pm 0.003}$ $0.350_{\pm 0.027}$ $0.003_{\pm 0.000}$
missMDA $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
VAEAC $0.004_{\pm 0.000}$ $0.380_{\pm 0.025}$ $0.003_{\pm 0.000}$
MIWAE $0.010_{\pm 0.001}$ $0.405_{\pm 0.034}$ $0.003_{\pm 0.000}$
not-MIWAE $0.006_{\pm 0.001}$ $0.410_{\pm 0.031}$ $0.003_{\pm 0.000}$
EGC $0.006_{\pm 0.000}$ $0.890_{\pm 0.023}$ $0.036_{\pm 0.002}$
MaCoDE $0.007_{\pm 0.001}$ $0.430_{\pm 0.040}$ $0.003_{\pm 0.000}$
Dataset kings
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.016_{\pm 0.002}$ $0.527_{\pm 0.033}$ $0.012_{\pm 0.000}$
GAIN $0.045_{\pm 0.005}$ $0.345_{\pm 0.026}$ $0.012_{\pm 0.000}$
missMDA $0.034_{\pm 0.003}$ $0.295_{\pm 0.028}$ $0.013_{\pm 0.000}$
VAEAC $0.009_{\pm 0.001}$ $0.636_{\pm 0.030}$ $0.012_{\pm 0.000}$
MIWAE $0.005_{\pm 0.000}$ $0.709_{\pm 0.018}$ $0.013_{\pm 0.000}$
not-MIWAE $0.008_{\pm 0.001}$ $0.645_{\pm 0.034}$ $0.013_{\pm 0.000}$
EGC $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
MaCoDE $0.007_{\pm 0.002}$ $0.809_{\pm 0.034}$ $0.013_{\pm 0.000}$
Dataset letter
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.021_{\pm 0.001}$ $0.412_{\pm 0.019}$ $0.014_{\pm 0.000}$
GAIN $0.043_{\pm 0.005}$ $0.394_{\pm 0.023}$ $0.013_{\pm 0.000}$
missMDA $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
VAEAC $0.015_{\pm 0.001}$ $0.512_{\pm 0.022}$ $0.014_{\pm 0.000}$
MIWAE $0.014_{\pm 0.003}$ $0.756_{\pm 0.032}$ $0.033_{\pm 0.006}$
not-MIWAE $0.021_{\pm 0.004}$ $0.738_{\pm 0.038}$ $0.025_{\pm 0.004}$
EGC $0.019_{\pm 0.002}$ $0.969_{\pm 0.010}$ $0.102_{\pm 0.010}$
MaCoDE $0.040_{\pm 0.003}$ $0.400_{\pm 0.019}$ $0.018_{\pm 0.003}$
Dataset loan
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE $0.009_{\pm 0.001}$ $0.740_{\pm 0.052}$ $0.027_{\pm 0.000}$
GAIN $0.026_{\pm 0.006}$ $0.480_{\pm 0.053}$ $0.027_{\pm 0.000}$
missMDA $0.016_{\pm 0.003}$ $0.600_{\pm 0.073}$ $0.031_{\pm 0.001}$
VAEAC $0.010_{\pm 0.001}$ $0.700_{\pm 0.033}$ $0.028_{\pm 0.000}$
MIWAE $0.005_{\pm 0.001}$ $0.940_{\pm 0.031}$ $0.029_{\pm 0.000}$
not-MIWAE $0.006_{\pm 0.001}$ $0.860_{\pm 0.043}$ $0.029_{\pm 0.000}$
EGC $-_{\pm-}$$-_{\pm-}$$-_{\pm-}$
MaCoDE $0.009_{\pm 0.003}$ $0.860_{\pm 0.060}$ $0.029_{\pm 0.001}$
@@
Table 16: Q3: Multiple imputation for each dataset under MNARL at 0.3 missingness. The means and standard errors of the mean across 10 repeated experiments are reported. Due to computational issues, the number of multiple imputations is set to 10 for the covtype dataset, while for other datasets it is set to 100. $\downarrow$ denotes lower is better.

### v2

@@ Appendix A Appendix / supplemental material
9:     end for
10:     $\hat{Q}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{Q}_{m}^{(s)}$
11:     $\hat{U}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{U}_{m}^{(s)}+(M+1)/M\cdot\frac{1}{M-1}\sum_{m=1}^{M}(\hat{U}_{m}^{(s)}-\bar{U}^{(s)})^{2}$, where $\bar{U}^{(s)}=\frac{1}{M}\sum_{m=1}^{M}\hat{U}_{m}^{(s)}$
12:  end for
13:  Bias: $\frac{1}{S}\sum_{s=1}^{S}|\hat{Q}^{(s)}-Q^{*}|$
14:  Coverage: $\frac{1}{S}\sum_{s=1}^{S}\mathbb{I}(Q^{*}\in(\hat{Q}^{(s)}\pm 1.96\cdot\sqrt{\hat{U}^{(s)}}))$
15:  Confidence interval length: $\frac{1}{S}\sum_{s=1}^{S}2\cdot 1.96\cdot\sqrt{\hat{U}^{(s)}}$
For Q3, we selected the following multiple imputation models that can handle heterogeneous tabular datasets: MICE (van Buuren and Groothuis-Oudshoorn 2011), GAIN (Yoon, Jordon, and van der Schaar 2018), missMDA (Josse and Husson 2016), VAEAC (Ivanov, Figurnov, and Vetrov 2019), MIWAE (Mattei and Frellsen 2019), not-MIWAE (Ipsen, Mattei, and Frellsen 2021), and EGC (Zhao, Townsend, and Udell 2022).
We assess the effectiveness of multiple imputations by employing interval inference for the population mean, which was proposed by Rubin (Rubin and Schenker 1986). However, since obtaining the population mean is not feasible, we utilize the sample column-wise mean of the complete dataset as a parameter of interest (Lee and Huber 2021; Zhang et al. 2023). The evaluation procedure for multiple imputations is outlined in Algorithm 2, and we utilize torch.random.manual_seed() to set seeds in performing multiple imputations. We report the mean and standard error of bias, coverage, and confidence interval length across 10 different random seeds, all continuous columns, and datasets (van Buuren 2012). Note that we do not split the data into training and test sets (Ipsen, Mattei, and Frellsen 2021).
Remark 3.
We acknowledge that existing imputation methods have been evaluated using RMSE (root mean square error), a metric for assessing single imputation methods (Ipsen, Mattei, and Frellsen 2021; Mattei and Frellsen 2019; Nazábal et al. 2020; Muzellec et al. 2020; Zhao et al. 2023a; Jarrett et al. 2022). However, since our objective focuses on distributional learning rather than recovering missing values, we adopt the evaluation procedure proposed in (Rubin and Schenker 1986; van Buuren 2012), better suited for assessing multiple imputation methods.
Following (Muzellec et al. 2020; Jarrett et al. 2022; Zhao et al. 2023a), we generate the missing value mask for each dataset with three mechanisms in four settings. (MCAR) In the MCAR setting, each value is masked according to the realization of a Bernoulli random variable with a fixed parameter. (MAR) In the MAR setting, for each experiment, a fixed subset of variables that cannot have missing values is sampled. Then, the remaining variables have missing values according to a logistic model with random weights, which takes the non-missing variables as inputs. A bias term is fitted using line search to attain the desired proportion of missing values. (MNAR) Finally, two different mechanisms are implemented in the MNAR setting. The first, MNARL, is identical to the previously described MAR mechanism, but the inputs of the logistic model are then masked by an MCAR mechanism. Hence, the logistic model’s outcome depends on missing values. The second mechanism, MNARQ, samples a subset of variables whose values in the lower and upper $p$th percentiles are masked according to a Bernoulli random variable, and the values in-between are left not missing.
Model Bias $\downarrow$ Coverage Width $\downarrow$
MICE .$010_{\pm.001}.$$845_{\pm.019}.$$\mathbf{.040}_{\pm.002}$
GAIN .$019_{\pm.002}.$$633_{\pm.033}.$$\mathbf{.040}_{\pm.002}$
missMDA .$015_{\pm.001}.$$700_{\pm.022}.$$043_{\pm.002}.$
VAEAC .$\underline{.008}_{\pm.001}$$905_{\pm.016}.$$\mathbf{.040}_{\pm.002}$
MIWAE .$\mathbf{.006}_{\pm.000}$$\underline{.952}_{\pm.012}$$043_{\pm.002}.$
not-MIWAE .$\mathbf{.006}_{\pm.000}$$\mathbf{.949}_{\pm.012}$$\underline{.042}_{\pm.002}$
EGC .$\mathbf{.006}_{\pm.000}$${.996}_{\pm.004}.$$058_{\pm.002}.$
MaCoDE(MAR) .$\mathbf{.006}_{\pm.000}$$963_{\pm.009}.$$051_{\pm.003}.$
Table 5: Q3: Multiple imputation under MAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better. Coverage close to 0.95 indicates better performance. The best value is bolded, and the second best is underlined.
Results
The missing data mechanism within the parentheses refers to the mechanism applied to the dataset on which MaCoDE was trained. Table 5 indicates that MaCoDE consistently exhibits competitive performance against all baseline models across metrics assessing multiple imputation performances, including bias, coverage, and confidence interval length. This suggests our proposed approach can support multiple imputations for deriving statistically valid inferences from missing data with the MAR mechanism.
Additional Experiments
Privacy Preservability
@@ Results
  • Evaluation metrics: The $k$-anonymity property (Sweeney 2002) is a measure used to assess the level of privacy protection in synthetic data. It ensures that each individual’s information in the dataset cannot be distinguished from that of at least $k-1$ other individuals. In other words, each record in the dataset has at least $k-1$ similar records in terms of quasi-identifiers, attributes that could potentially identify a subject. A higher $k$ value implies a higher level of anonymity and better privacy preservation.
@@
    DCR (Distance to Closest Record) (Park et al. 2018; Zhao et al. 2021) is defined as the distances between all real training and synthetic samples. A higher DCR value indicates more effective privacy preservation, indicating a lack of overlap between the real training data and the synthetic samples. Conversely, an excessively large DCR score suggests a lower quality of the generated synthetic dataset. Therefore, the DCR metric provides insights into both the privacy-preserving capability and the quality of the synthetic dataset.
@@
    Attribute disclosure (Choi et al. 2017; Matwin et al. 2015) refers to the situation where attackers can uncover additional covariates of a record by leveraging a subset of covariates they already possess, along with similar records from the synthetic dataset. To quantify the extent to which attackers can accurately identify these additional covariates, we employ classification metrics. Higher attribute disclosure metrics indicate an increased risk of privacy leakage, implying that attackers can precisely infer unknown variables. In terms of privacy concerns, attribute disclosure can be considered a more significant issue than membership inference attacks, as attackers are assumed to have access to only a subset of covariates for a given record.
@@
  • Evaluation procedure: We evaluate the $k$-anonymity following the approach described in (Qian, Davis, and van der Schaar 2023)  [^14] https://github.com/vanderschaarlab/synthcity. Additionally, similar to (Zhao et al. 2021), we define DCR as the $5^{th}$ percentile of the $L_{2}$ distances between all real training samples and synthetic samples. Since DCR relies on $L_{2}$ distance and continuous variables, it is computed solely using continuous variables. We assess attribute disclosure using the methodology outlined in (Choi et al. 2017).
@@
Figure 4: Trade-off between privacy and quality. Left: feature selection performance (synthetic data quality). Right: DCR (privacy preservability). The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. Error bars represent the standard errors of the mean.
@@
Table 6: Trade-off between privacy and quality (statistical fidelity and machine learning utility). The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. $\uparrow$ ($\downarrow$) denotes higher (lower) is better.
@@
Detailed Experimental Results
Q1. Synthetic Data Quality
@@
Q1. Visualization of Marginal Histogram
@@
Q2: Synthetic Data Quality in Scenarios with Incomplete Training Dataset
@@
Q3: Multiple Imputation Performance
@@
MICE $0.014_{\pm 0.001}$ $0.795_{\pm 0.022}$ $0.040_{\pm 0.002}$
GAIN $0.024_{\pm 0.002}$ $0.591_{\pm 0.037}$ $0.040_{\pm 0.002}$
missMDA $0.018_{\pm 0.001}$ $0.644_{\pm 0.015}$ $0.045_{\pm 0.002}$
VAEAC $0.010_{\pm 0.001}$ $0.874_{\pm 0.022}$ $0.041_{\pm 0.002}$
MIWAE $0.008_{\pm 0.000}$ $0.936_{\pm 0.014}$ $0.045_{\pm 0.002}$
notMIWAE $0.008_{\pm 0.000}$ $0.910_{\pm 0.013}$ $0.044_{\pm 0.002}$
EGC $0.005_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.060_{\pm 0.002}$
MaCoDE(MCAR) $0.008_{\pm 0.000}$ $0.950_{\pm 0.011}$ $0.053_{\pm 0.002}$
@@
MICE $0.010_{\pm 0.001}$ $0.845_{\pm 0.019}$ $0.040_{\pm 0.002}$
GAIN $0.019_{\pm 0.002}$ $0.633_{\pm 0.033}$ $0.040_{\pm 0.002}$
missMDA $0.015_{\pm 0.001}$ $0.700_{\pm 0.022}$ $0.043_{\pm 0.002}$
VAEAC $0.008_{\pm 0.001}$ $0.905_{\pm 0.016}$ $0.040_{\pm 0.002}$
MIWAE $0.006_{\pm 0.000}$ $0.952_{\pm 0.012}$ $0.043_{\pm 0.002}$
notMIWAE $0.006_{\pm 0.000}$ $0.949_{\pm 0.012}$ $0.042_{\pm 0.002}$
EGC $0.006_{\pm 0.000}$ $0.996_{\pm 0.004}$ $0.058_{\pm 0.002}$
MaCoDE(MAR) $0.006_{\pm 0.000}$ $0.963_{\pm 0.009}$ $0.051_{\pm 0.003}$
@@
MICE $0.012_{\pm 0.001}$ $0.790_{\pm 0.026}$ $0.037_{\pm 0.002}$
GAIN $0.026_{\pm 0.002}$ $0.538_{\pm 0.033}$ $0.040_{\pm 0.002}$
missMDA $0.020_{\pm 0.001}$ $0.592_{\pm 0.025}$ $0.045_{\pm 0.002}$
VAEAC $0.011_{\pm 0.001}$ $0.851_{\pm 0.023}$ $0.041_{\pm 0.002}$
MIWAE $0.009_{\pm 0.001}$ $0.920_{\pm 0.015}$ $0.045_{\pm 0.002}$
notMIWAE $0.009_{\pm 0.001}$ $0.908_{\pm 0.012}$ $0.044_{\pm 0.002}$
EGC $0.007_{\pm 0.000}$ $0.994_{\pm 0.004}$ $0.061_{\pm 0.002}$
MaCoDE(MNARL) $0.008_{\pm 0.000}$ $0.948_{\pm 0.013}$ $0.052_{\pm 0.003}$
@@
MICE $0.007_{\pm 0.001}$ $0.880_{\pm 0.018}$ $0.040_{\pm 0.002}$
GAIN $0.011_{\pm 0.001}$ $0.819_{\pm 0.020}$ $0.040_{\pm 0.002}$
missMDA $0.011_{\pm 0.001}$ $0.785_{\pm 0.016}$ $0.043_{\pm 0.002}$
VAEAC $0.006_{\pm 0.001}$ $0.909_{\pm 0.020}$ $0.040_{\pm 0.002}$
MIWAE $0.005_{\pm 0.000}$ $0.969_{\pm 0.009}$ $0.043_{\pm 0.002}$
notMIWAE $0.005_{\pm 0.000}$ $0.958_{\pm 0.010}$ $0.043_{\pm 0.002}$
EGC $0.004_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.054_{\pm 0.002}$
MaCoDE(MNARQ) $0.005_{\pm 0.000}$ $0.944_{\pm 0.012}$ $0.051_{\pm 0.003}$
Table 12: Q3: Multiple imputation under MCAR, MAR, MNARL, and MNARQ at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes the lower is better.
@@
Table 13: Q3: Multiple imputation for each dataset under MCAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better.
@@
Table 14: Q3: Multiple imputation for each dataset under MAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better. $\downarrow$ denotes lower is better.